whats 8 times the difference between a number and 5 is equal to the number plus 2. whats the number?
Answer by rjd41(9)

8(x-5)=x+2
8x-40=x+2
8x-x=7x 2+40=42
7x=42
x=6
check:
6-5=1, 8(1)=8
6+2=8
the number is 6